A 10,000 watt radio station transmits at 535 kHz.
Calculate the energy of a single photon at the transmitted frequency
1
Expert's answer
2013-03-12T09:04:44-0400
E=hv, E-energy, h-Planck constant, v-frequency
E=6.63×10^(-34)*535*10^3=3.55*10^(-28) (J) Photon energydoesn't depend on the power of radio station